# Entanglement in quenched extended Su-Schrieffer-Heeger model with anomalous dynamical quantum phase transitions

### v1

@@ I Introduction
The formal establishment of the theory of dynamical quantum phase transitions (DQPTs) in 2013 [1], followed by the studies of its fundamental properties [2, 3, 4, 5, 6] and successful experimental realization [7, 8, 9, 10], leads to an alternative approach in studying non-equilibrium physics in quantum many-body systems. DQPTs have been witnessed and substantially studied in spin models [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], topological models [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] and lattice gauge theory [58, 59, 60, 61, 62], via various protocols such as a quench or a periodical drive [63, 64, 65, 66, 67, 68, 69, 70] to bring the system out of equilibrium. There are three types of DQPTs. The type-I DQPTs (DPTs-I) considers the kinks at the transition point in the long-time average of the order parameter of a system [5, 6, 71, 72, 73]. On the other hand, the type-II DQPTs (DPTs-II), which will be the main focus of this work, concerns the transient time dynamics, and are defined by the zeros of the Loschmidt amplitude (LA), the overlap of the time-evolved state driven by the unitary evolution of the final Hamiltonian $H_{f}$ onto the initial state |$\ket{\psi_{0}^{i}}$
@@
Table 1: Summary of the features of the critical momentum $k^{*}$, critical time $t^{*}$, the dynamical vector $\vec{r}_{k}$, entanglement entropy $S_{A}$, correlation matrix spectrum $\{\xi_{j}\}$ evolution in different dynamical phases. The regular and anomalous phase refer to the case where the number of critical momenta is equal to and larger than the winding number difference, respectively (see the text for the details). “N.A.” in the table stands for “not applicable” since there is no DQPT.
@@
In topological models, it is found that there can be more than one critical momenta (and correspondingly critical time scales) [43, 53, 54, 55, 56, 57]. These critical momenta originate from the different topological structures between the initial and final phases, in which their number must be at least the difference in the winding number between the two phases for 1D models. The cases with excessive critical momenta are referred to as accidental in Ref. [53]. For 2D topological models, robust DQPTs are present if there are at least one node in the $k$-space wavefunction overlap, and such nodes are guaranteed when the pre- and postquenched phases have a different Chern number [55, 56]. The 2D case in general has areas of zeros of the Loschmidt amplitude distributed on the complex time plane and a continuum of solutions of the critical times, but it is found in a number of cases that have line-like distribution of zeros whose critical time usually agrees well with the Loschmidt rate non-analytic peak [80].
In this work, we show that the previously referred to accidental DQPTs in topological models in fact contains three notions, distinguished by quenches (1) within the same phase, (2) across the phase boundary but without a change in the winding numbers, and (3) across the boundary with a change in the winding numbers. We demonstrate our findings via an extended version of the Su-Schrieffer-Heeger (SSH) model, a commonly studied topological model in condensed matter physics, due to its simplicity and accessibility to experiments [81]. The extended SSH model possesses a rich equilibrium phase diagram, allowing us to probe the quantum dynamics of quenches involving different topological and the trivial phases. Specifically, we realize for quenches between trivial and the topological phases, and between phases with the same winding number, there exists regions in the dynamical phase diagram, where the number of critical momenta $n_{k^{*}}>|\Delta W|$, the absolute difference of winding number between the two phases. Distinctive features are found between the cases of $n_{k^{*}}=|\Delta W|$ and $n_{k^{*}}>|\Delta W|$. For quench between phases with the same winding number, DQPTs are not just found to be possible for quench within the same phase, but also across the equilibrium phase boundary. We reckon the observations may be a generic nonequilibrium feature for topological models beyond nearest-neighbor hoppings, and further formulate and term these kinds of accidental DQPTs the type-B anomalous DQPTs. Note that the type-B anomalous DQPTs here is different from the anomalous DQPTs (we call them type-A anomalous DQPTs for unambiguity) described in Ref. [5, 72, 83, 84, 82, 6, 16, 85, 86], which emphasizes the minimum in rate functions before the first critical time.
@@ II Dynamical Phase Diagram of the extended SSH model
It was shown that the trajectory of $\vec{r}_{k}(t)$ in the first Brillouin zone at the first critical time reveals distinct features for quenches from the equilibrium phase boundary [25]. Here, we find the trajectories of $\vec{r}_{k}$ exhibit different characteristics for the quenches involving the regular and anomalous regions in the dynamical phase diagram, and the results will be presented below.
@@
Figure 4: Locations of $k^{*}$’s in the complex plane in the vicinity of the anomalous dynamical phase boundaries of an $L=400$ system for quenches (a) $t_{c}^{i}=-0.5,t_{d}^{i}=1$ to $t_{c}^{f}$ ranging from 1 to 1.6 with $t_{d}^{f}=-0.7$ and (b) $t_{c}^{i}=-1.2,t_{d}^{i}=-1.8$ to $t_{d}^{f}$ ranging from $-1.4$ to $-0.8$ with $t_{c}^{f}=-0.2$. Color dots from dim to dark colors represent increasing final parameters.
@@
Figure 2 shows the dynamical phase diagram defined by the number of critical momenta $n_{k^{*}}$ for an initial state taken from each of the four equilibrium phases. Interestingly, for the quench between $W_{0}$ and $W_{1}$ phases, a dynamical phase boundary separating the $1k^{*}$ (the regular) and $3k^{*}$ (the anomalous) regions within the same equilibrium phase is observed (Fig. 2(a) and 2(b)). We also find that DQPTs can occur for quenches without crossing an equilibrium phase boundary. Especially for quenches within the $W_{2}$ or $W_{-1}$ phases (Fig. 2(c) and 2(d)), there exists a finite region in the dynamical phase diagram that DQPTs take place within the same phase. These observations motivate us to define the type-B anomalous DQPT as the regions where $n_{k*}$ exceeds $|\Delta W|$. In later parts we will elucidate theoretically that all these type-B anomalous DQPTs originate from the same cause.
@@
Here, we define the regions with $n_{k^{*}}$ exceeding $|\Delta W|$ as the anomalous regions and present the possible cause and ways to distinguish them from regular DQPTs of the type-B anomalous DQPTs. Note that in both anomalous cases presented in Table 1, two additional critical momenta are found (comparing column 1 & 3, and column 2 & 4). The two extra $k^{*}$’s are originated from a complex conjugate pair of $k^{*}$ in $\vec{d}_{i}(k^{*})\cdot\vec{d}_{f}(k^{*})=0$, where they share the same real part. As the postquenched parameters enter the anomalous region (the $3k^{*}$ region for quench between $W_{0}$ and $W_{1}$, and the $2k^{*}$ region for quench within the $W_{-1}$ or $W_{2}$ phases), the imaginary part of the two critical momenta vanishes and they transform to the real critical momenta, giving rise to extra real critical times. This ensures that the parity of $n_{k^{*}}$ is conserved in both regions. For example, a $W_{2}\rightarrow W_{2}$ quench gives either zero $k^{*}$ or 2 $k^{*}$’s, while a $W_{1}\rightarrow W_{0}$ quench gives either 1 $k^{*}$ or 3 $k^{*}$’s.
@@
The emergence of the complex-conjugate pair is visualized in Fig. 4, in which it illustrates the locations of $k^{*}$’s on the complex plane around the dynamical phase boundary. The complex conjugate pairs zone in the real axis when the final quench parameters approach the dynamical phase boundaries. The two critical momenta overlap right at the dynamical phase boundary, implying only one additional critical time resulted. Notice however that this is to be differentiated from the degenerate critical modes observed in quenched clustered Ising model, which is related to the structure of its energy spectrum [38]. When quenched further into the anomalous regions, the two critical momenta separate farther apart from each other and two distinct critical times emerge.
@@
Apart from the location of $k^{*}$’s, the vector $\vec{r}_{k}$ also reveals distinct features between regular and type-B anomalous regions, as summarized in Table 1. Figure 5 shows the trajectories of $\vec{r}_{k}$ for the respective quench cases in Fig. 3 at the first critical time. For the regular case where $n_{k^{*}}=|\Delta W|$, each $k^{*}$ is “assigned” a separate loop in the trajectory of $\vec{r}_{k}$. It is clearly seen in Fig. 5(a), quenching from $W_{1}$ phase to $1k^{*}$ region in $W_{0}$ phase where we have one loop for one critical momentum, and Fig. 5(c), quenching from the bottom $W_{2}$ phase to the left $W_{-1}$ phase in the phase diagram shown where we have three loops, each containing one $k^{*}$. On the other hand, for the anomalous case with quench to $3k^{*}$ in Fig. 5(b) and the quench within the bottom $W_{2}$ phase in Fig. 5(d), the critical momentum pair gathers in the same loop instead of staying in two separate loops, which further suggests the two extra critical momenta have the same origin. One thing to notice is the extra loop for quench $W_{1}\rightarrow W_{0}$ to $1k^{*}$ region in Fig. 5(a) that has no critical momentum contained in it. This extra loop corresponds to the imaginary critical momentum pair that are not yet landed onto the real axis and cosine range. For the postquenched Hamiltonian approaching the dynamical boundary, the loop will eventually touch the origin, marking the emergence of the $k^{*}$ pair. This feature still holds true for the SSH model with even longer range hoppings as discussed in Appendix A.
@@ III.0.1 EE in the original SSH model
We start from the original SSH model in equilibrium. Two premises are stated below in order to facilitate discussions: 1. The $W_{0}$ phase favors stronger $t_{a}$ hoppings ($\eta>0$) while the $W_{1}$ phase favors stronger $t_{b}$ hoppings ($\eta<0$) [95, 94]; and 2. physically, EE reflects the information lack on the whole system when one only considers a certain subsystem. In other words, higher entropy corresponds to a higher lack of knowledge in the subsystem and vice versa. Combining the above statements, we argue that the equilibrium EE of $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ (refer to Fig. 1(a) & (b)) should be larger in the $W_{1}{\color[rgb]{0,0,0}/W_{0}}$ phase than in the $W_{0}{\color[rgb]{0,0,0}/W_{1}}$ phase. On one hand, the system favors $t_{a}{\color[rgb]{0,0,0}/t_{b}}$ over $t_{b}{\color[rgb]{0,0,0}/t_{a}}$ hoppings in the $W_{0}{\color[rgb]{0,0,0}/W_{1}}$ phase. On the other hand, $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ contains a $t_{a}{\color[rgb]{0,0,0}/t_{b}}$ bond but not a $t_{b}{\color[rgb]{0,0,0}/t_{a}}$ bond. Therefore, $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ contains more information about the $W_{0}{\color[rgb]{0,0,0}/W_{1}}$ phase, resulting in a smaller value of EE at equilibrium (see Fig. 6(b) $t=0$). Similarly, the $W_{1}{\color[rgb]{0,0,0}/W_{0}}$ phase favors $t_{b}{\color[rgb]{0,0,0}/t_{a}}$ over $t_{a}{\color[rgb]{0,0,0}/t_{b}}$ hoppings, resulting in a higher EE for the subsystem $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ due to its lack of information (see Fig. 6(a) $t=0$). These are especially true for the blue curves, as they demonstrate quenches deeper inside the corresponding phases. Therefore, in the original SSH model, the EE can be large or small in both the $W_{0}$ and the $W_{1}$ phases depending on the choice of the subsystem.
@@
Now we move on to the discussion of the behavior near DQPT for the original SSH model. Here, our main focus is on the EE of small subsystems. For the half-block EE, there is an interesting discussion in Ref. [44], where the authors demonstrated the relation between the time evolution of the two-point correlations and the oscillatory behaviour in EE. In Fig. 6(a), a local minimum in EE for subsystem $L_{A}=1$ is attained around DQPT for quench $W_{1}\rightarrow W_{0}$ phase (red and blue curves). Following the discussion above, this implies an influx of information from the surroundings to the subsystem $L_{A}=1$ in the vicinity of DQPT, which is induced by the quench. Figure 6(b) shows quenches where EEs achieve local maxima near DQPT. The maximal behavior of EE suggests the outflow of information from the corresponding subsystem to the surroundings in the vicinity of DQPT. One can also notice the exact opposite happened to the subsystem $L_{B}=1$, where the EE is maximum and minimum for $W_{1}\rightarrow W_{0}$ and $W_{0}\rightarrow W_{1}$, respectively, around the critical time. We can understand this as the ”role” of $W_{0}$ and $W_{1}$ can be switched by choosing an appropriate subsystem.

### v2

@@ I Introduction
The formal establishment of the theory of dynamical quantum phase transitions (DQPTs) in 2013 [1], followed by the studies of its fundamental properties [2, 3, 4, 5, 6] and successful experimental realization [7, 8, 9, 10], leads to an alternative approach in studying non-equilibrium physics in quantum many-body systems. DQPTs have been witnessed and substantially studied in spin models [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 98], topological models [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 99] and high energy physics [58, 59, 60, 61, 62, 100, 101], via various protocols such as a quench or a periodical drive [63, 64, 65, 66, 67, 68, 69, 70] to bring the system out of equilibrium. There are three types of DQPTs. The type-I DQPTs (DPTs-I) considers the kinks at the transition point in the long-time average of the order parameter of a system [5, 6, 71, 72, 73]. On the other hand, the type-II DQPTs (DPTs-II), which will be the main focus of this work, concerns the transient time dynamics, and are defined by the zeros of the Loschmidt amplitude (LA), the overlap of the time-evolved state driven by the unitary evolution of the final Hamiltonian $H_{f}$ onto the initial state |$\ket{\psi_{0}^{i}}$
@@
Table 1: Summary of the features of the critical momentum $k^{*}$, critical time $t^{*}$, the dynamical vector $\vec{r}_{k}$, entanglement entropy $S_{A}$, correlation matrix spectrum $\{\xi_{j}\}$ evolution in different dynamical phases. The regular and anomalous phase refer to the case where the number of critical momenta is equal to and larger than the absolute winding number difference, respectively (see the text for the details). “N.A.” in the table stands for “not applicable” since there is no DQPT.
@@
In topological models, it is found that there can be more than one critical momenta (and correspondingly critical time scales) [43, 53, 54, 55, 56, 57]. These critical momenta originate from the different topological structures between the initial and final phases, in which their number must be at least the difference in the winding number between the two phases for 1D models. The cases with excessive critical momenta are referred to as accidental cases in Ref. [53]. For 2D topological models, robust DQPTs are present if there are at least one node in the $k$-space wavefunction overlap, and such nodes are guaranteed when the pre- and postquenched phases have a different Chern number [55, 56]. The 2D case in general has areas of zeros of the Loschmidt amplitude distributed on the complex time plane and a continuum of solutions of the critical times, but it is found in a number of cases that have line-like distribution of zeros whose critical time usually agrees well with the Loschmidt rate non-analytic peak [80].
In this work, we show that the previously referred to accidental DQPTs in topological models in fact contains three notions, distinguished by quenches (1) within the same phase, (2) across the phase boundary but without a change in the winding numbers, and (3) across the boundary with a change in the winding numbers. We demonstrate our findings via an extended version of the Su-Schrieffer-Heeger (SSH) model, a commonly studied topological model in condensed matter physics, due to its simplicity and accessibility to experiments [81]. The extended SSH model possesses a rich equilibrium phase diagram, allowing us to probe the quantum dynamics of quenches involving different topological and the trivial phases. Specifically, we realize for quenches between trivial and the topological phases, and between phases with the same winding number, there exists regions in the dynamical phase diagram, where the number of critical momenta $n_{k^{*}}>|\Delta W|$, the absolute difference of winding number between the two phases. Distinctive features are found between the cases of $n_{k^{*}}=|\Delta W|$ and $n_{k^{*}}>|\Delta W|$. For quench between phases with the same winding number, DQPTs are not just found to be possible for quench within the same phase, but also across the equilibrium phase boundary. We reckon the observations may be a generic nonequilibrium feature for topological models beyond nearest-neighbor hoppings, and further formulate and term these kinds of accidental DQPTs the type-B anomalous DQPTs. Note that the type-B anomalous DQPTs here is different from the anomalous DQPTs (we call them type-A anomalous DQPTs for unambiguity) described in Ref. [5, 72, 83, 84, 82, 6, 16, 85, 86, 100], which emphasizes the minimum in rate functions before the first critical time.
@@ II Dynamical Phase Diagram of the extended SSH model
It was shown that the trajectory of $\vec{r}_{k}$ in the first Brillouin zone at the first critical time reveals distinct features for quenches from the equilibrium phase boundary [25]. Here, we find the trajectories of $\vec{r}_{k}$ exhibit different characteristics for the quenches involving the regular and anomalous regions in the dynamical phase diagram, and the results will be presented below.
@@
Figure 4: Locations of $k^{*}$’s in the complex plane in the vicinity of the anomalous dynamical phase boundaries of an $L=400$ system for quenches (a) $t_{c}^{i}=-0.5,t_{d}^{i}=1$ to $t_{c}^{f}$ ranging from 1 to 1.6 with $t_{d}^{f}=-0.7$ and (b) $t_{c}^{i}=-1.2,t_{d}^{i}=-1.8$ to $t_{d}^{f}$ ranging from $-1.4$ to $-0.8$ with $t_{c}^{f}=-0.2$. Color dots from light to dark colors represent increasing final parameters.
@@
Figure 2 shows the dynamical phase diagram defined by the number of critical momenta $n_{k^{*}}$ for an initial state taken from each of the four equilibrium phases. Interestingly, for the quench between $W_{0}$ and $W_{1}$ phases, a dynamical phase boundary separating the $1k^{*}$ (the regular) and $3k^{*}$ (the anomalous) regions within the same equilibrium phase is observed (Figs. 2(a) and 2(b)). We also find that DQPTs can occur for quenches without crossing an equilibrium phase boundary. Especially for quenches within the $W_{2}$ or $W_{-1}$ phases (Figs. 2(c) and 2(d)), there exists a finite region in the dynamical phase diagram that DQPTs take place within the same phase. These observations motivate us to define the type-B anomalous DQPT as the regions where $n_{k^{*}}$ exceeds $|\Delta W|$. In later parts we will elucidate theoretically that all these type-B anomalous DQPTs originate from the same cause.
@@
Here, we define the regions with $n_{k^{*}}$ exceeding $|\Delta W|$ as the type-B anomalous regions and present the possible cause and ways to distinguish them from regular DQPTs. Note that in both anomalous cases presented in Table 1, two additional critical momenta are found (comparing column 1 and 3, and column 2 and 4). The two extra $k^{*}$’s are originated from a complex conjugate pair of $k^{*}$ in $\vec{d}_{i}(k^{*})\cdot\vec{d}_{f}(k^{*})=0$, where they share the same real part. As the postquenched parameters enter the anomalous region (the $3k^{*}$ region for quench between $W_{0}$ and $W_{1}$, and the $2k^{*}$ region for quench within the $W_{-1}$ or $W_{2}$ phases), the imaginary part of the two critical momenta vanishes and they transform to the real critical momenta, giving rise to extra real critical times. This ensures that the parity of $n_{k^{*}}$ is conserved in both regions. For example, a $W_{2}\rightarrow W_{2}$ quench gives either zero $k^{*}$ or 2 $k^{*}$’s, while a $W_{1}\rightarrow W_{0}$ quench gives either 1 $k^{*}$ or 3 $k^{*}$’s.
@@
The emergence of the complex-conjugate pair is visualized in Fig. 4, in which it illustrates the locations of $k^{*}$’s on the complex plane around the dynamical phase boundary. The complex conjugate pairs zone in the real axis when the final quench parameters approach the dynamical phase boundaries. The two critical momenta overlap right at the dynamical phase boundary, implying only one additional critical time resulted. Notice, however, that this is to be differentiated from the degenerate critical modes observed in quenched clustered Ising model, which is related to the structure of its energy spectrum [38]. When quenched further into the anomalous regions, the two critical momenta separate farther apart from each other on the real axis and two distinct critical times emerge.
@@
Apart from the location of $k^{*}$’s, the vector $\vec{r}_{k}$ also reveals distinct features between regular and type-B anomalous regions, as summarized in Table 1. Figure 5 shows the trajectories of $\vec{r}_{k}$ for the respective quench cases in Fig. 3 at the first critical time. For the regular case where $n_{k^{*}}=|\Delta W|$, each $k^{*}$ is “assigned” a separate loop in the trajectory of $\vec{r}_{k}$. It is clearly seen in Fig. 5(a), quenching from $W_{1}$ phase to $1k^{*}$ region in $W_{0}$ phase where we have one loop for one critical momentum, and Fig. 5(c), quenching from the bottom $W_{2}$ phase to the left $W_{-1}$ phase in the phase diagram shown where we have three loops, each containing one $k^{*}$. On the other hand, for the anomalous case with quench to $3k^{*}$ in Fig. 5(b) and the quench within the bottom $W_{2}$ phase in Fig. 5(d), the critical momentum pair gathers in the same loop instead of staying in two separate loops, which further suggests the same origin of the two extra critical momenta. This feature still holds true for the SSH model with even longer range hoppings as discussed in Appendix A. One thing to notice is the extra loop for quench $W_{1}\rightarrow W_{0}$ to $1k^{*}$ region in Fig. 5(a) that has no critical momentum contained in it. This extra loop corresponds to the imaginary critical momentum pair that are not yet landed onto the real axis and cosine range. For the postquenched Hamiltonian approaching the dynamical boundary, the loop will eventually touch the origin, marking the emergence of the $k^{*}$ pair.
@@ III.0.1 EE in the original SSH model
We start from the original SSH model in equilibrium. Two premises are stated below in order to facilitate discussions: 1. The $W_{0}$ phase favors stronger $t_{a}$ hoppings ($\eta>0$) while the $W_{1}$ phase favors stronger $t_{b}$ hoppings ($\eta<0$) [95, 94]; and 2. physically, EE reflects the information lack on the whole system when one only considers a certain subsystem. In other words, higher entropy corresponds to a higher lack of knowledge in the subsystem and vice versa. Combining the above statements, we argue that the equilibrium EE of $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ (refer to Figs. 1(a) and 1(b)) should be larger in the $W_{1}{\color[rgb]{0,0,0}/W_{0}}$ phase than in the $W_{0}{\color[rgb]{0,0,0}/W_{1}}$ phase. On one hand, the system favors $t_{a}{\color[rgb]{0,0,0}/t_{b}}$ over $t_{b}{\color[rgb]{0,0,0}/t_{a}}$ hoppings in the $W_{0}{\color[rgb]{0,0,0}/W_{1}}$ phase. On the other hand, $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ contains a $t_{a}{\color[rgb]{0,0,0}/t_{b}}$ bond but not a $t_{b}{\color[rgb]{0,0,0}/t_{a}}$ bond. Therefore, $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ contains more information about the $W_{0}{\color[rgb]{0,0,0}/W_{1}}$ phase, resulting in a smaller value of EE at equilibrium (see Fig. 6(b) $t=0$). Similarly, the $W_{1}{\color[rgb]{0,0,0}/W_{0}}$ phase favors $t_{b}{\color[rgb]{0,0,0}/t_{a}}$ over $t_{a}{\color[rgb]{0,0,0}/t_{b}}$ hoppings, resulting in a higher EE for the subsystem $L_{A}=1{\color[rgb]{0,0,0}/L_{B}=1}$ due to its lack of information (see Fig. 6(a) $t=0$). These are especially true for the blue curves, as they demonstrate quenches deeper inside the corresponding phases. Therefore, in the original SSH model, the EE can be large or small in both the $W_{0}$ and the $W_{1}$ phases depending on the choice of the subsystem.
@@
Now we move on to the discussion of the behavior near DQPT for the original SSH model. Here, our main focus is on the EE of small subsystems. For the half-block EE, there is an interesting discussion in Ref. [44], where the authors demonstrated the relation between the time evolution of the two-point correlations and the oscillatory behaviour in EE. In Fig. 6(a), a local minimum in EE for subsystem $L_{A}=1$ is attained around DQPT for quench $W_{1}\rightarrow W_{0}$ phase (orange and light blue curves). Following the discussion above, this implies an influx of information from the surroundings to the subsystem $L_{A}=1$ in the vicinity of DQPT induced by the quench. Figure 6(b) shows quenches where EEs achieve local maxima near DQPT. The maximal behavior of EE suggests an outflow of information from the corresponding subsystem to the surroundings in the vicinity of DQPT. One can also notice the exact opposite happened to the subsystem $L_{B}=1$, where the EE is maximum and minimum for $W_{1}\rightarrow W_{0}$ and $W_{0}\rightarrow W_{1}$, respectively, around the critical time. We can understand this as the ”role” of $W_{0}$ and $W_{1}$ can be switched by choosing an appropriate subsystem.